OpenStudy (winsteria):

Every pay period, Steven budgets 7/12 of his paycheck for expenses, 1/6 for savings, and the rest for recreation. His paycheck is $864.80. How much money does Steven put aside for recreation each pay period? CAN SOMEONE PLEASE HELP ME WITH THIS ONE I'LL GIVE YOU A MEDAL AND A FAN???

OpenStudy (dangerousjesse):

Common denominator is twelve, so \(\large\frac{7}{12}+\frac{2}{12}=\frac{9}{12}=\frac{3}{4}\)

OpenStudy (dangerousjesse):

\(\large\frac{3}{4}=.75\)

OpenStudy (dangerousjesse):

Multiply 864.8 by .25 to find the answer.

OpenStudy (winsteria):

so it's $216.2 yea???